OpenStudy (anonymous):

Help please??? Find S12 for the series 1 + 1.5 + 2.25 + 3.375 +...

OpenStudy (anonymous):

r=1.5/1=1.5=3/2\[sn=\frac{ a(r^n-1) }{ r-1 }\] n=12 calculate

OpenStudy (anonymous):

is it 257.5?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

\[1\frac{ \left( \frac{ 3 }{ 2 } \right)^{12}-1 }{ \frac{ 3 }{ 2 }-1 }\] \[=\frac{ \frac{ 531441 }{ 4096 }-1 }{ \frac{ 1 }{ 2 } }=2\left( \frac{ 531441-4096 }{4096 } \right)\] \[=\frac{ 527345 }{ 2048 }=257.49 \] so you are absolutely correct.